### 05 Dec 2023 v1

# Approximate Nash Equilibria Algorithms for Network Design Games

Abstract
We consider a weighted network design game, where selfish players chooses paths in a network to minimize their cost. The cost function of each edge in the network is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$. We first show the existence of $\alpha$-approximate Nash equilibrium and prove the upper bound of $\alpha$ is $O(log_{2}(W))$, where $W$ is the sum of the weight of all players. Furthermore, considering that compute the $\alpha$-approximate Nash equilibrium is PLS-complete, we assume that $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables on $[0,1]$. In this case, we show that $\epsilon$-better response dynamics can compute the $\alpha$-approximate Nash Equilibrium in polynomial time by proving the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players and the number of edges in the network.
Keywords:
Approximate Nash equilibrium Affine linear cost function Better response dynamics
1 Introduction
Along with the development of many networks including the Internet, the number of agents using networks is increasing, all of whom are motivated by self-interest and optimize an individual objective function. This naturally leads us to study the behavior of these non-cooperative agents in the game. Moreover, the outcome of the interactions of selfish agents is also worthy of analysis, and the stable outcomes are called Nash equilibrium. Furthermore, the study of Nash equilibrium gives rise to a number of new issues. Specifically, since the overall performance of the system brought about by the selfish agents making their own decisions may be worse than that brought about by centralized authority decision-making, we need an indicator to measure the degree to which the global performance of the system has deteriorated. Papadimitriou [1] first proposed using the term price of anarchy to refer to the deterioration of system performance. Moreover, there is more than one Nash equilibrium in some types of games, but the price of anarchy only studies the deterioration in the worst Nash equilibrium. Then Anshelevich et al. [3] introduced the term price of stability to refer to the ratio of the cost of best Nash equilibrium to the optimum network cost.
On the other hand, computing a Nash equilibrium is essential when the game is given from a computational perspective. However, Fabrikant et al. [5] proved that computing a Nash equilibrium of a congestion game is PLS-complete. Furthermore, Ackermann et al. [6] showed that the PLS-completeness holds in network design game and other kinds of games. After that, scholars have turned their direction to computing approximate Nash equilibrium. However, Skopalik et al. [7] proved that computing an $\alpha$-PNE is PLS-complete for any given constant $\alpha$. Therefore, it is necessary to further restrict the game ensure that there is a method to compute the (approximate) Nash equilibrium in polynomial time.
1.1 Related Works
1.1.1 Network Design Game
Papadimitriou [1] first proposed using algorithmic ideas and game theory to analyze the cost of network. Anshelevich et al. [3] proposed the network design with Shapley cost sharing, which uses a cost-sharing mechanism. However Anshelevich et al. [3] only proved few results when we players are weighted. H. Chen and T. Roughgarden [8] showed that weighted Shapley network design game does not admit a Nash equilibrium. Moreover, they proved the existence of $\alpha$-approximate Nash equilibrium and gives a upper bound of $\alpha$. However, no previous study has been found for the case where the cost function is a linear function to the total weight of the players that choose the edge.
1.2 Our contributions
Our model extend the cost function in the model proposed by Anshelevich et al. [3], which is a constant $c_{e}$ for each edge $e$ in the network. In our model, the cost function of each edge in the network is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$. We show the existence of $\alpha$-approximate Nash equilibrium and prove the upper bound of $\alpha$ is $O(log_{2}(W))$, where $W$ is the sum of the weight of all players.
Furthermore, as Skopalik et al. [7] proved that computing an $\alpha$-PNE is PLS-complete for any given constant $\alpha$ in congestion games, our model assume that $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables on $[0,1]$, which means that the cumulative distribution function of $a_{e},b_{e}$ is upper bounded by a constant $\phi\geq 1$. In this case, we show that $\epsilon$-better response dynamics can compute the $\alpha$-approximate Nash Equilibrium in polynomial time by proving the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi$, the number of players $n$ and the number of edges $m$ in the network.
2 Model and Preliminaries
We represent an arbitrary weighted network design game $\Gamma$ by tuple $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ with components defined below.
  • P1: $G=(V,E)$ is a directed graph, where $V$ is the vertex set of $G$ and $E$ is the edge set of $G$. Here we further define $m=|E|$, which is the number of edges in graph $G$.
  • P2: $\mathcal{N}$ is a collection of $n$ different players. Moreover, each player $i\in\mathcal{N}$ is identified with a source-sink vertex pair $(s_{i},t_{i})$.
  • P3: $\mathcal{P}=(\mathcal{P}_{i})_{i\in\mathcal{N}}$, where $\mathcal{P}_{i}$ denote the set of simple $s_{i}-t_{i}$ paths. Furthermore, each player $i$ chooses a path $P_{i}\in\mathcal{P}_{i}$ from its source to its sink.
  • P4: $\mathcal{C}=(c_{e})_{e\in E}$, where $c_{e}$ is the non-negative cost of the edge $e$.
  • P5: $\mathcal{W}=(w_{i})_{i\in\mathcal{N}}$, where each component $w_{i}>0$ is a positive weight of player $i\in\mathcal{N}$. When $w_{i}=w_{j}$ for any $i,j\in\mathcal{N}$, then $\Gamma$ is called unweighted. Otherwise, $\Gamma$ is called weighted.
All players are selfish, and each of them chooses a path independently. Then it results in a pure strategy profile $P=(P_{i})_{i\in\mathcal{N}}$, where $P_{i}$ denotes the simple $s_{i}-t_{i}$ path used by player $i\in\mathcal{N}$. For a given profile and an edge $e\in E$, we define $U_{e}(P):=\{i|i\in\mathcal{N},e\in P_{i}\}$ as the “user” of edge $e$. Moreover, we define $W_{e}=\sum_{i\in U_{e}(P)}w_{i}$ as the total weight of the players that choose a path containing edge $e$. Furthermore, for a given profile and a player $i$, the cost share $c^{i}_{e}$ of an edge $e\in E$ is $c_{e}\cdot w_{i}/W_{e}$. Then the cost to player $i$ in an profile is the sum of its cost shares: $c_{i}(P)=\sum_{e\in P_{i}}c^{i}_{e}$. In addition, the cost $c(P)$ of a profile $P$ is defined as $\sum_{i\in\mathcal{N}}c_{i}(P)$, which can be also written as $\sum_{e\in\cup_{i\in\mathcal{N}}P_{i}}c_{e}$.
To state our discussion precisely, for each player $i\in\mathcal{N}$, we denote by $P_{-i}$ a sub-profile $(P_{i}^{\prime})_{i^{\prime}\in\mathcal{N}\backslash\{i\}}$ of strategies used by “opponents” $i^{\prime}\in\mathcal{N}\backslash\{i\}$ of player $i$, and use $P=(P_{i},P_{-i})$ when we need to show explicitly the strategy $P_{i}$ used by player $i$
Definition 1
(Pure Strategy Nash Equilibrium) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ a pure strategy Nash equilibrium(PNE) if, for each $i$, $P_{i}$ minimizes $c_{i}(P)$ over all paths in $\mathcal{P}_{i}$ while keeping $P_{j}$ fixed for $j\neq i$. That is:
$c_{i}(P)=c_{i}(P_{i},P_{-i})\leq c_{i}(P^{\prime}_{i},P_{-i})$ (1)
for any $i\in\mathcal{N}$, any $P^{\prime}_{i}\in\mathcal{P}_{i}$.
In fact, Inequality (1) states that $P_{i}$ is a best-response to $P_{-i}$ for player $i$. Therefore, there is no incentive for any player $i\in\mathcal{N}$ to unilaterally change strategy when the profile is already a PNE.
To prove PNE exists, a common approach is to prove that there is a potential function in the game. Here we give the definition of the potential function.
Definition 2
(Exact Potential Function) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an (exact) potential function if
$\Phi(P^{\prime}_{i},P_{-i})-\Phi(P_{i},P_{-i})=c_{i}(P^{\prime}_{i},P_{-i})-c_{i}(P_{i},P_{-i})$ (2)
for any $i\in\mathcal{N}$, any $P_{i},P^{\prime}_{i}\in\mathcal{P}_{i}$, any $P_{-i}\in\prod_{j\in\mathcal{N}\backslash\{i\}}\mathcal{P}_{j}$.
If there is a potential function in a finite game, since the potential function quantifies the cost reduction of a unilateral change of strategy exactly, then the global minimums of the potential function are PNEs of the game.
However, in some cases, the game may not admit a PNE. Then we have to consider a relaxed notion of $\alpha$-approximate pure strategy Nash equilibrium instead.
Definition 3
($\alpha$-Approximate Pure Strategy Nash Equilibrium) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game and an arbitrary constant $\alpha\geq 1$. We call a profile $P=(P_{i})_{i\in\mathcal{N}}$ an $\alpha$-approximate pure strategy Nash equilibrium ($\alpha$-APNE) if
$c_{i}(P)=c_{i}(P_{i},P_{-i})\leq\alpha\cdot c_{i}(P^{\prime}_{i},P_{-i})$ (3)
for any $i\in\mathcal{N}$, any $P^{\prime}_{i}\in\mathcal{P}_{i}$.
In fact, Inequality (3) states that a unilateral change of strategy reduces the cost at a rate of at most $\frac{c_{i}(P_{i},P_{-i})-c_{i}(P^{\prime}_{i},P_{-i})}{c_{i}(P_{i},P_{-i})}\leq 1-\frac{1}{\alpha}$ in an $\alpha$-APNE. Moreover, we call a deviation $\alpha$-improving if the deviation decreases the cost incurred by the player by at least an $\alpha$ multiplicative factor. Therefore, $\alpha$-APNE are profiles from which no $\alpha$-improving deviations exist. We further point out that the smaller the constant $\alpha$ is, the closer the profile $P$ would approximate a PNE.
Correspondingly, we can prove the existence of $\alpha$-APNE by finding the approximate potential function for the game.
Definition 4
($\alpha$-Approximate Potential Function) Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. For any player $i\in\mathcal{N}$, $P_{i},P^{\prime}_{i}\in\mathcal{P}_{i}$ satisfies $P_{i}\rightarrow P^{\prime}_{i}$ is an $\alpha$-improving deviation. We call a real-valued function $\Phi:\prod_{i\in\mathcal{N}}\mathcal{P}_{i}\rightarrow[0,\infty)$ an $\alpha$-approximate potential function if
$\Phi(P^{\prime}_{i},P_{-i})-\Phi(P_{i},P_{-i})\leq\alpha c_{i}(P^{\prime}_{i},P_{-i})-c_{i}(P_{i},P_{-i})$ (4)
for any $P_{-i}\in\prod_{j\in\mathcal{N}\backslash\{i\}}\mathcal{P}_{j}$.
If there is a $\alpha$-approximate potential function in a finite game, since the potential function strictly decreases in every $\alpha$-improving deviation, then there will be no $\alpha$-improving deviation after a finite number of steps, which means the game admits an $\alpha$-APNE.
Without loss of generality, we assume $\min_{i\in\mathcal{N}}w_{i}=1$ in the following discussion, and we define $w_{max}=\max_{i\in\mathcal{N}}w_{i}$ and $W=\sum_{i\in\mathcal{N}}w_{i}$ to simplify our discussion.
$\phi$-smooth random variable. The probability density function of each random variable is up to $\phi$($\phi\geq 1$).
3 Existence of $\alpha$-APNE
Anshelevich et al. [3] proved that the unweighted network design game admits a PNE, and pointed out that weighted Shapley network design game does not admit a PNE. Therefore, it is essential to turn the attention to $\alpha$-APNE.
Theorem 3.1
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game. If the cost function is affine linear, namely $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $a_{e},b_{e}>0$ are only related to the edge $e$ and $W_{e}$ is the total weight of the players that choose a path containing edge $e$, then the game admits an $\alpha$-APNE, where $\alpha=O(log_{2}(W))$.
Before proving the theorem above, we first present a lemma.
Lemma 1
Let $x,y\in[1,\infty)$, then we have:
  1. 1.
    $log_{2}(1+x+y)-log_{2}(1+x)\geq\frac{y}{x+y}$;
  2. 2.
    $log_{2}(1+x+y)-log_{2}(1+x)<log_{2}[e(1+y)]\cdot\frac{y}{x+y}$.
Proof
We first point out $(1+\frac{1}{x})^{x}$ increases monotonically and $\lim_{x\to\infty}(1+\frac{1}{x})^{x}=e$. For part 1, if $y\geq x+1$, then
$log_{2}(1+x+y)-log_{2}(1+x)=log_{2}(1+\frac{y}{x+1})\geq 1\geq\frac{y}{x+y}$
if $y<x+1$, then we have
$(1+\frac{y}{x+1})^{\frac{x+y}{y}}\geq(1+\frac{y}{x+1})^{\frac{x+1}{y}}\geq 2$
Therefore, we have $log_{2}(1+\frac{y}{x+1})\geq\frac{y}{x+y}$, which completes the proof for part 1.
For part 2, we have
$(1+\frac{y}{x+1})^{\frac{x+y}{y}}$ $=(1+\frac{y}{x+1})^{\frac{x+1}{y}}(1+\frac{y}{x+1})^{\frac{y-1}{y}}$
$<(1+\frac{y}{x+1})^{\frac{x+1}{y}}(1+\frac{y}{x+1})$
$\leq e(1+y)$
Similarly to the first part, $log_{2}(1+\frac{y}{x+1})\leq log_{2}[e(1+y)]\cdot\frac{y}{x+y}$, which completes the proof for part 2.
With the Lemma 1, we now prove Theorem 3.1.
Proof
For a weighted network design game, We define a function $\Phi$ as follows:
$\Phi(P)=\sum_{e\in E}c_{e}(W_{e})log_{2}(1+W_{e})$
Claim
$\Phi$ is an $\alpha$-approximate potential function for the weighted network design game.
To prove the claim, we need to show that $\epsilon$-improving deviations strictly decrease the function $\Phi$. Specifically, we consider an $\epsilon$-improving deviation of player $i$ from the profile $P=(P_{i},P^{\prime}_{-i})$ to the profile $P^{\prime}=(P^{\prime}_{i},P_{-i})$. Without loss of generality, we can assume that $P_{i}$ and $P^{\prime}_{i}$ are disjoint. Otherwise, we can replace $P_{i}$ and $P^{\prime}_{i}$ with $P_{i}\backslash P^{\prime}_{i}$ and $P^{\prime}_{i}\backslash P_{i}$ in the following discussion. By the definition of $\alpha$-improving, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{c_{e}(W_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}\leq\sum_{e\in P_{i}}{c_{e}(W_{e})}\cdot\frac{w_{i}}{W_{e}}.$
Here, $W_{e}$ denotes the total weight on edge $e$ before the deviation. Since $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{(a_{e}(W_{e}+w_{i})+b_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}\leq\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\cdot\frac{w_{i}}{W_{e}}.$ (5)
Then we can derive the following:
$\Delta\Phi$ $=\sum_{e\in P^{\prime}_{i}}{(a_{e}W_{e}+b_{e})}(log_{2}(1+W_{e}+w_{i})-log_{2}(1+W_{e}))$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}(log_{2}(1+W_{e})-log_{2}(1+W_{e}-w_{i}))$
$\,\quad+\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})log_{2}(1+W_{e}+w_{i})-\sum_{e\in P_{i}}(a_{e}w_{i})log_{2}(1+W_{e}-w_{i})$ (6)
$<\sum_{e\in P^{\prime}_{i}}{(a_{e}W_{e}+b_{e})}log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}}$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\frac{w_{i}}{W_{e}}+\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})log_{2}(1+W_{e}+w_{i})$ (7)
$=\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})(log_{2}(1+W_{e}+w_{i})+log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}})$
$\,\quad+\sum_{e\in P^{\prime}_{i}}b_{e}log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\frac{w_{i}}{W_{e}}$ (8)
$<\alpha\sum_{e\in P^{\prime}_{i}}{(a_{e}(W_{e}+w_{i})+b_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\cdot\frac{w_{i}}{W_{e}}$ (9)
$=\alpha c_{i}(P^{\prime})-c_{i}(P),$ (10)
where $\alpha=2log_{2}(1+W+w_{max})$. Therefore, we have proved
$\Phi(P^{\prime})-\Phi(P)<\alpha c_{i}(P^{\prime})-c_{i}(P),$ (11)
which completes the proof of the claim. Moreover, since there are a finite number of profiles, the theorem can be directly derived from the claim.
4 Better Response Dynamics
Initially, we propose an algorithm to find the $\alpha$-APNE of the game $\Gamma$ as followings:
Algorithm 1 $\epsilon$-Approximate Better Response Dynamics($\epsilon$-ABRD)
Input:
A weighted network design game $\Gamma=(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$; a strategy profile $P\in\mathcal{P}$
Output:
An $\alpha$-APNE of game $\Gamma$
  1. 1.
    While $P$ is not an $\alpha$-APNE do
  2. 2.
    Choose $i\in\mathcal{N}$, $P^{\prime}_{i}\in\mathcal{P}_{i}$ such that $(1+\epsilon)\alpha c_{i}(P^{\prime}_{i},P_{-i})<c_{i}(P)=c_{i}(P_{i},P_{-i})$
  3. 3.
    $P\leftarrow(P^{\prime}_{i},P_{-i})$
  4. 4.
    end while
  5. 5.
    return $P$
Since computing $\alpha$-APNE is PLS-complete, we need to make further assumption on the game.
Definition 5
Let $X$ be a random variable, we say $X$ is a $\phi$-smooth random variable if and only if $X$ is defined on $[0,1]$ and the cumulative distribution function of $X$ is upper bounded by a constant $\phi\geq 1$.
Before proving the algorithm above can fine the $\alpha$-APNE in a polynomial time, we first prove two lemmas as preparation.
Lemma 2
Let $X_{1},X_{2},...,X_{n};Y_{1},Y_{2},...,Y_{n}$ be $\phi$-smooth independent random variables on [0,1]. Let $g(x)=\min\{\frac{\alpha}{x}ln\frac{\alpha}{x},n^{\lambda}\}$. Then for any $\lambda\in N$ and $\alpha\geq 1$, we have
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]\leq\alpha\phi n\cdot ln(\alpha\phi n)\cdot(\lambda+1)^{2}$
Proof
For two random variables $S,T$, we define $S\preceq T$ if and only if $F_{T}(x)\leq F_{S}(x)$ for any $x\in\mathbb{R}$, where $F_{S}(x),F_{T}(x)$ are the cumulative distribution function of $S,T$ respectively. To prove the lemma, we define two sets of independent uniformly distributed random variables $S_{1},S_{2},...,S_{n};T_{1},T_{2},...,T_{n}$ on $[0,\frac{1}{\phi}]$.
Claim
For any $i\in\{1,2,...,n\}$, we have
$min_{i\in\{1,2,...,n\}}S_{i}\preceq min_{i\in\{1,2,...,n\}}X_{i},min_{i\in\{1,2,...,n\}}T_{i}\preceq min_{i\in\{1,2,...,n\}}Y_{i}.$
To prove the claim, we firstly prove $F_{X_{i}}(t)\leq F_{S_{i}}(t)$ for any $t\in\mathbb{R}$, then we can derive $F_{Y_{i}}(t)\leq F_{T_{i}}(t)$ similarly.
Since $S_{i}$ is defined on $[0,\frac{1}{\phi}]$, we have $F_{S_{i}}(t)=1$ for any $t\geq\frac{1}{\phi}$, then we have $F_{X_{i}}(t)\leq F_{S_{i}}(t)$ definitely. When $0\leq t\leq\frac{1}{\phi}$, since $X_{i}$ is a $\phi$-smooth random variable for any $i\in\{1,2,...,n\}$, its probability density function is upper-bounded by $\phi$. Then we have
[…]
for all $0\leq t\leq\frac{1}{\phi}$. Therefore, we have $F_{X_{i}}(t)\leq F_{S_{i}}(t)$ and $F_{Y_{i}}(t)\leq F_{T_{i}}(t)$. Furthermore, since $X_{1},X_{2},...,X_{n}$ are independent random variables, we have
$F_{min_{i\in\{1,2,...,n\}}X_{i}}(t)$ $=1-\prod^{n}_{i=1}(1-F_{X_{i}}(t))$
$\leq 1-\prod^{n}_{i=1}(1-F_{S_{i}}(t))$
$=F_{min_{i\in\{1,2,...,n\}}S_{i}}(t),$
which completes the proof of the claim.
Moreover, we can find that $g$ is nonincreasing with respect to $x$. From the Proposition 9.1.2 in [4], we have
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]\leq\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{S_{i}+T_{i}\})].$ (12)
We define $W=\frac{1}{\min_{i\in\{1,2,...n\}}\{S_{i}+T_{i}\}}$, then $W\in[\frac{\phi}{2},\infty)$. Since $S_{i},T_{i}$ are independent uniformly distributed random variables on $[0,\frac{1}{\phi}]$, we have
$F_{W}(w)=Prob[W\leq w]=\prod^{n}_{i=1}Prob[S_{i}+T_{i}\geq\frac{1}{w}]=(1-\frac{\phi^{2}}{2w^{2}})^{n}.$
For $i=0,1,...,\lambda$, we define $Q_{i}$ as the random event that $W\in[\frac{\phi}{2}\cdot n^{i},\frac{\phi}{2}\cdot n^{i+1})$. Moreover, we let $Q_{\lambda+1}$ denote the random event that $W\in[\frac{\phi}{2}\cdot n^{\lambda+1},\infty)$. Then, we can upper bound the expectation in (12) by:
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]$ $\leq\mathbb{E}[\min\{\alpha\cdot Wln(\alpha\cdot W),n^{\lambda}\}]$
$\leq\sum^{\lambda}_{i=0}Prob[Q_{i}]\cdot\mathbb{E}[\alpha\cdot Wln(\alpha\cdot W)]+Prob[Q_{\lambda+1}]\cdot n^{\lambda}$
$\leq\sum^{\lambda}_{i=0}[F_{W}(\frac{\phi}{2}\cdot n^{i+1})-F_{W}(\frac{\phi}{2}\cdot n^{i})]\cdot\frac{\alpha\phi}{2}\cdot n^{i+1}\cdot ln(\frac{\alpha\phi}{2}\cdot n^{i+1})$
$\,\quad+[1-F_{W}(\frac{\phi}{2}\cdot n^{\lambda+1})]\cdot n^{\lambda}$
$\leq\frac{\alpha\phi}{2}\cdot ln(\frac{\alpha\phi n}{2})\cdot\sum^{\lambda}_{i=0}[F_{W}(\frac{\phi}{2}\cdot n^{i+1})-F_{W}(\frac{\phi}{2}\cdot n^{i})]\cdot n^{i+1}\cdot(i+1)$
$\,\quad+[1-F_{W}(\frac{\phi}{2}n^{\lambda+1})]\cdot n^{\lambda}$
$=\frac{\alpha\phi}{2}\cdot ln(\frac{\alpha\phi n}{2})\cdot\sum^{\lambda}_{i=0}[(1-\frac{2}{n^{2i+2}})^{n}-(1-\frac{2}{n^{2i}})^{n}]\cdot n^{i+1}\cdot(i+1)$
$\,\quad+[1-(1-\frac{2}{n^{2\lambda+2}})^{n}]\cdot n^{\lambda}$
$\leq\alpha\phi n\cdot ln(\alpha\phi n)\cdot(\lambda+1)^{2}.$
Therefore, we complete the proof of the lemma.
Lemma 3
Let $x\in(0,1)$, then we have:
$-\frac{1}{ln(1-x)}\leq\frac{1}{x}$
Proof
Since $x\in(0,1)$, we have $-ln(1-x)\geq 0$. Moreover, it is easy to find out that $ln(1-x)=ln(\frac{1}{1-x})$ and $x=1-\frac{1}{\frac{1}{1-x}}$. By defining $t=\frac{1}{1-x}$, we only need to prove
$lnt\geq 1-\frac{1}{t}$
for any $t>1$. We define $f(t)=lnt+\frac{1}{t}-1$, then $f^{\prime}(t)=\frac{t-1}{t^{2}}\geq 0$. Therefore, $f(t)\geq f(1)=0$, which completes the proof.
Now we begin to prove that $\epsilon$-ABRD always find the $\alpha$-APNE in a polynomial time. It is easy to see that we only need to prove the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi,n,m$.
Theorem 4.1
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game and $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables. Then, the expected number of iterations of $\epsilon$-ABRD is at most $O(\frac{1}{\epsilon}W^{3}\phi n^{3}m^{3}ln(W\phi nm))$.
Proof
We recall that the approximate potential function $\Phi$ of the game $\Gamma$ is
$\Phi(P)=\sum_{e\in E}c_{e}(W_{e})log_{2}(1+W_{e}),$
which can be upper-bounded by
$\Phi(P)\leq\sum_{e\in E}c_{e}(W)log_{2}(1+W_{e})\leq m\max_{e\in E}\{a_{e}W+b_{e}\}log_{2}(1+W).$ (13)
To simplify our discussion, we define $c_{max}=\max_{e\in E}\{a_{e}W+b_{e}\}$. For the lower bound of the approximate potential function $\Phi$, we first propose a fact as following:
$log_{2}(1+x)+log_{2}(1+y)\geq log_{2}(1+x+y)$
for any $x,y>0$. And since $\min_{i\in\mathcal{N}}w_{i}=1$ and each player uses at least one edge, then $\Phi$ can be lower bounded by
$\Phi(P)\geq\sum_{e\in E}\min_{e\in E}\{a_{e}+b_{e}\}log_{2}(1+W_{e})\geq\min_{e\in E}\{a_{e}+b_{e}\}log_{2}(1+W).$ (14)
To simplify our discussion, we define $c_{min}=\min_{e\in E}\{a_{e}+b_{e}\}$.
Furthermore, according to algorithm 1, we assume that we move $T$ steps $P^{0}\rightarrow P^{1}\rightarrow...\rightarrow P^{T}$ during the execution of our $\epsilon$-ABRD and $P\rightarrow P^{\prime}$ is one of them. Since $c_{i}(P)\geq\frac{c_{min}}{W}$, we can derive the following from Inequality (11) and Inequality (13):
$\Phi(P)-\Phi(P^{\prime})$ $>c_{i}(P)-\alpha c_{i}(P^{\prime})$
$>\frac{\epsilon}{1+\epsilon}c_{i}(P)$
$\geq\frac{\epsilon}{(1+\epsilon)W}c_{min}$
$\geq\frac{\epsilon c_{min}}{(1+\epsilon)\cdot W\cdot log_{2}(1+W)\cdot m\cdot c_{max}}\Phi(P).$ (15)
So we have
$\Phi(P^{\prime})<(1-\epsilon^{\prime})\Phi(P),\quad\text{where}\;\epsilon^{\prime}=\frac{\epsilon c_{min}}{(1+\epsilon)Wlog_{2}(1+W)mc_{max}}.$ (16)
Since Inequality (16) holds for any move during the execution of the $\epsilon$-ABRD, we can derive the following from Inequality (13) and Inequality (14):
$c_{min}log_{2}(1+W)\leq\Phi(P^{T})<(1-\epsilon^{\prime})^{T}\Phi(P^{0})\leq(1-\epsilon^{\prime})^{T}mc_{max}log_{2}(1+W).$
Therefore, we take logarithms at both sides of the inequality and get
$Tln(1-\epsilon^{\prime})>ln\frac{c_{min}}{mc_{max}}.$ (17)
Since $10<\epsilon^{\prime}<$ 1, we can derive the following upper bound on the number of steps for our $\epsilon$-ABRD from Lemma 3 and Inequality (17):
$T$ $<-\frac{1}{ln(1-\epsilon^{\prime})}ln\frac{mc_{max}}{c_{min}}$
$\leq\frac{1}{\epsilon^{\prime}}ln\frac{mc_{max}}{c_{min}}$
$=(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}}.$ (18)
On the other hand, we recall that the approximate potential function $\Phi$ is strictly decreasing at every step of the $\epsilon$-ABRD, which means no profile appears more than once during the $\epsilon$-ABRD. Therefore, the total number of iterations cannot be larger than the number of different profiles, which is up to $(2^{n})^{m}=2^{nm}$. Combining this with Inequality (4) we have
$T\leq\min\{(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}},2^{nm}\}.$ (19)
Since $c_{min}=\min_{e\in E}\{a_{e}+b_{e}\}$ and $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables on $[0,1]$, then we can derive the following from Lemma 2:
$\mathbb{E}[T]$ $\leq(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\mathbb{E}[\min\{\frac{m(1+W)}{c_{min}}ln\frac{m(1+W)}{c_{min}},n^{nmlog_{n}2}\}]$
$\leq(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\cdot m(1+W)\phi\cdot n\cdot ln(m(1+W)\phi\cdot n)(nmlog_{n}2+1)^{2}$
$\leq 4\cdot(1+\frac{1}{\epsilon})W^{3}\cdot\phi\cdot n^{3}\cdot m^{3}\cdot ln(W\cdot\phi\cdot n\cdot m)$
$=O(\frac{1}{\epsilon}W^{3}\cdot\phi\cdot n^{3}\cdot m^{3}\cdot ln(W\cdot\phi\cdot n\cdot m)),$
which completes the proof of the theorem.